what is the standard form of the equation of the line through the given point and slope
through:(3,-1), slope=-1

Answer by Alan3354(69443)

what is the standard form of the equation of the line through the given point and slope
through:(3,-1), slope=-1
----------------
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.
------
then convert to standard form.
============
You'll benefit from the practice. I won't.